# The Class 0 protostars in Orion: Characterizing the properties of their magnetized envelopes

### v1

@@ Abstract
We present a study connecting the physical properties of protostellar envelopes to the morphology of the envelope-scale magnetic field. We use the ALMA polarization observations of 55 young prtostars at 0.87 mm on $\sim 400-3000$ au scales from the B-field Orion Protostellar Survey (BOPS) to infer the envelope-scale magnetic field and both dust and gas emission on comparable scales to measure the envelope properties. We find that the protostellar envelopes with compact polarized dust emission tend to have lower envelope masses, than the sources with more extended envelopes. We also find that protostars showing hourglass-field morphologies tend to have lower velocity dispersions in their envelopes, whereas systems with spiral-field morphologies have higher velocity dispersion. Combining with the disk properties taken from the Orion VLA/ALMA Nascent Disk and Multiplicity (VANDAM) survey, we connect envelope properties to fragmentation. Our results suggest that envelope mass may not correlate with fragmentation, whereas turbulence appears to promote fragmentation. On the other hand, we find that fragmentation is suppressed in systems with pinched magnetic fields, suggesting that the magnetic field play a role on providing additional support against gravitational collapse, and the formation of an hourglass-like field may coincide with enhanced magnetic braking that removes angular momentum and hinders the formation of embedded disks. Nevertheless, significant misalignment between magnetic field and outflow axes tends to reduce magnetic braking, leading to the formation of larger disks.
@@ 1 Introduction
Fragmentation is thought to be a critical process in the formation of star and stellar systems (e.g., Matsumoto et al., 2003). Initially, turbulence induces chaotic motions and density fluctuations across various scales, compressing the gas in certain regions and leading to localized density enhancements (e.g., Mac Low et al., 2004). Within these dense regions, gravitational forces overcome thermal pressure if the mass exceeds the Jeans mass. As the cloud collapses, large-structure fragment into smaller, dense cores, each of which form individual stars or multiple star systems (e.g., McKee et al., 2007). This hierarchical fragmentation continues until individual cores reach densities high enough to resist further fragmentation (e.g., Beuther et al., 2015). However, the specific mechanisms by which a cloud core fragments to ultimately form a multiple system or small cluster remain uncertain.
@@
Theoretical and numerical studies (e.g., Maury et al., 2022; Offner et al., 2023) suggest that cores with some angular momentum are prone to fragmentation into multiple objects, indicating the need for additional mechanisms to reduce the fragmentation to the observed levels. Since non-ideal MHD effects do not inhibit protostellar multiplicity (e.g., Wurster et al., 2019), further research points to turbulence and magnetic fields (hereafter, B-fields) as playing a pivotal role in the fragmentation process (e.g., Padoan et al., 2001; Price et al., 2009). For instance, compressive turbulence promotes fragmentation and the formation of structures with high density contrasts, whereas solenoidal turbulence (e.g., Federrath et al., 2010; Girichidis et al., 2011) suppresses fragmentation and maintains a smoother density structure within the cloud. On the other hand, the ratio of rotational to magnetic energy has been found to influence core fragmentation (e.g., Machida et al., 2008). Moreover, B-field tends to suppress fragmentation by providing additional support against gravitational collapse, and/or through magnetic braking, which carries away the angular momentum from the inner regions of the core (e.g., Commerçon et al., 2011; Guszejnov et al., 2021; Mignon-Risse et al., 2021).
@@
Dust polarization is the most commonly used tracer of B-fields in star-forming regions, as “radiative torques” (e.g., Hoang et al., 2009; Andersson et al., 2015) tend to align spinning and elongated dust grains with their long axes perpendicular to the ambient field direction. Over the last decades, dust polarization observations carried out with (sub)millimeter interferometers have increasingly proven effective in mapping B-fields at the scales of cores ($\sim 10^{4}$ au) and envelopes ($\sim 10^{2}$ to $10^{4}$ au) (e.g. Girart et al., 1999, 2013; Cox et al., 2018; Galametz et al., 2018; Hull et al., 2019; Le Gouellec et al., 2020; Cortés et al., 2021). Theory predicts that B-field regulates the collapse of clouds, while ambipolar diffusion allows the formation of “supercritical” dense cores, where the gravitational force dominates over the magnetic support. As cloud cores collapse, field lines are dragged inward to be drawn into pinched morphology. This has been observed in various studies (e.g., Girart et al., 2006, 2009; Qiu et al., 2014; Beltrán et al., 2019; Le Gouellec et al., 2019; Hull et al., 2020). In the case of flux freezing, the pinched geometry in the pseudo-disk strengthens the B-field and carries away angular momentum, leading to catastrophic magnetic braking and potentially preventing the disk formation (e.g., Galli et al., 2006). However, disk may form if the magnetic and rotational axes are misaligned (e.g., Joos et al., 2012; Li et al., 2013; Machida et al., 2020). To statistically test these scenarios related to fragmentation and star/disk formation, it is crucial to understand the connection between envelope properties and the resulting fragmentation.
@@
Recently, the first results from the B-fields Orion protostellar survey (BOPS, with a resolution of $\sim 1^{\prime\prime}$, corresponding to $\sim$ 400 au, Huang et al., 2024) presented the largest polarization study of B-fields in protostellar envelopes using dust polarization observations of 61 protostars in Orion. We successfully mapped the B-fields morphology at scales ranging from several hundred to a few thousand au, and classified the protostars into several different morphological categories: standard hourglass (with the main axis coincident with the outflow axis), rotated hourglass (similar to the standard hourglass but rotated by $\sim 90^{\circ}$, with its axis perpendicular to the outflow), spiral configuration, and complex B-field structure. In addition, observations from a major Orion interferometric study, known as the Orion VLA/ALMA Nascent Disk and Multiplicity (VANDAM, Tobin et al., 2020), have identified protostellar disks in the Orion molecular clouds with a resolution of $\sim 0\farcs 1$ (corresponding to $\sim$ 40 au). This study has provided data on the radius and mass of these disks. Since all BOPS protostars were originally observed at high resolution in the VANDAM survey, the combination of these datasets enables a comprehensive test of the magnetic braking model and allows for a more effective linkage of magnetic and turbulent properties to the characteristics of protostellar disks and multiplicity.
@@
This paper builds upon the first results of BOPS (Huang et al., 2024) and is organized as follows: Section 2 introduces the observations and data reduction processes. Section 3 presents the methods for calculation of physical parameters, the velocity dispersion, the angle dispersion of the B-field, and the fragmentation level. A detailed discussion of the correlation between these parameters is provided in Section 4. Finally, the conclusions are summarized in Section 5.
@@ 2 Observations
The 870 $\mu$m (Band 7) dust polarization observations were taken with ALMA in 2019 November, and the spatial resolution and the Largest Angular Scale (LAS) are $\sim$1″and $\sim$8″, corresponding to a linear resolution of $\sim$400 and $\sim$3200 au at a distance of $\sim$400 pc (see Huang et al. 2024 for further details on the observation setup).
@@ 3 Analysis
Of the original 61 protostars presented in Huang et al. (2024), we excluded 6 objects due to a lack of data on either their polarization (HOPS-87S and ONC1N-8-S) or the embedded star/disk (e.g., insufficient information to determine their bolometric luminosity and embedded disk properties, ONC1N-4-5-EN, ONC1N-4-5-ES, ONC1N-6-7, and ONC1N-8-N). Additionally, 16 objects exhibited polarization measurements consistent with self-scattering in their disks and no B-field detections in the envelopes. These 16 objects had uniform polarization morphologies that aligned with the disk minor axes, a tell-tale sign of self-scattering (Kataoka et al., 2015, 2017; Stephens et al., 2017). Furthermore, 7 protostellar envelopes exhibited unresolved polarization measurements. This leaves us with a sample of 32 protostars with robust B-field measurements in their envelopes and complementary star and disk information. Among these remaining protostars, 6 exhibit standard hourglass field structures, 11 show rotated hourglass field morphologies, 9 have spiral magnetic fields, and 6 have complex or disordered field morphologies. The following sections presents the estimation of physical envelope parameters, velocity dispersion, the angle dispersion of B-field, and fragmentation level. These properties will then be compared to the disk properties to understand which factors could potentially influence the envelope fragmentation and whether the initial misalignment between the B-field and outflows could result in reduced magnetic braking during disk formation.
@@ 3.1 Physical Envelope Properties
where $d$ is the distance to the protostar as obtained from Huang et al. (2024), $\nu$ is the frequency corresponding to the observed wavelength, $B_{\nu}(T_{\rm dust})$ is the Planck function at the dust temperature of $T_{\rm dust}$, and $\kappa_{\nu}$ is the dust opacity. We adopt $\kappa_{870\mu{\rm m}}\approx 1.84\ {\rm cm}^{2}\ {\rm g}^{-1}$, which is appropriate for dusty particles with ice mantles at densities of approximately $10^{6}$ cm-3, as reported by Ossenkopf et al. (1994).
@@
Name $L_{\rm{bol}}$ $S_{\nu}$ $\overline{T}_{\rm env}$ $T_{\rm c}$ $M_{\rm env}$ $N_{\rm gas}$ $n_{\rm gas}$ $\sigma_{\rm nth}$ $\delta\phi$ $N_{F}$ $M_{\rm disk}$bbfootnotemark: $R_{\rm disk}$bbfootnotemark: $\Delta\theta$ccfootnotemark:
@@
HOPS-11 9.00 0.40 26.1 16.7 0.26 1.3E+23 5.4E+06 0.72 / 1.00 135 41.9 14.0
HOPS-87aafootnotemark: 36.49 3.89 37.0 17.6 1.46 7.3E+23 3.0E+07 0.70 17.6$\pm$1.4 1.00 628 124.1 49.1
HOPS-359 10.00 0.91 26.8 16.1 0.62 3.1E+23 1.3E+07 0.65 29.2$\pm$2.3 1.00 90 62.0 15.6
HOPS-395 0.50 0.58 20.2 11.7 0.49 2.4E+23 1.0E+07 0.70 15.0$\pm$2.0 2.00 24 14.2 0.1
HOPS-400 2.94 1.18 21.5 13.7 1.01 5.0E+23 2.1E+07 0.82 18.2$\pm$1.9 2.00 107 48.7 14.3
HOPS-407 0.71 0.61 20.2 13.7 0.59 2.9E+23 1.2E+07 0.64 10.5$\pm$1.0 1.00 83 64.1 16.5
@@
HH270IRS 7.70 0.49 25.1 16.6 0.35 1.7E+23 7.2E+06 0.77 33.8$\pm$4.1 2.00 161 92.3 68.3
HOPS-78 8.93 0.81 26.0 19.8 0.48 2.4E+23 9.9E+06 0.77 40.2$\pm$4.3 2.25 188 147.1 86.0
HOPS-168 48.07 0.51 39.6 18.1 0.17 8.4E+22 3.5E+06 0.90 28.4$\pm$3.9 2.00 65 65.8 63.4
HOPS-169 3.91 0.67 22.2 15.7 0.47 2.3E+23 9.7E+06 0.63 32.9$\pm$3.9 1.00 119 53.8 51.3
HOPS-288 135.47 1.60 51.4 18.8 0.43 2.1E+23 8.9E+06 1.38 24.5$\pm$2.9 3.00 263 144.3 80.0
HOPS-310 13.83 0.58 29.1 17.7 0.33 1.6E+23 6.8E+06 0.95 / 1.00 263 101.3 89.9
HOPS-317Saafootnotemark: 4.90 2.33 22.9 14.2 1.95 9.7E+23 4.0E+07 0.64 / 1.00 1059 269.1 83.4
HOPS-325 6.20 0.28 23.9 22.5 0.22 1.1E+23 4.6E+06 0.65 / 1.00 59 34.9 51.0
HOPS-370 360.86 1.08 65.6 25.4 0.21 1.0E+23 4.3E+06 1.22 33.8$\pm$3.1 1.00 114 112.8 86.8
HOPS-401 0.61 0.37 20.2 14.2 0.37 1.9E+23 7.7E+06 0.63 / 1.00 186 74.7 78.8
HOPS-409 8.18 0.38 25.4 22.5 0.23 1.2E+23 4.8E+06 0.78 17.0$\pm$2.5 1.00 111 112.1 73.3
@@
HOPS-96 6.19 0.44 23.8 19.9 0.29 1.4E+23 6.0E+06 0.57 / 1.00 135 41.4 52.4
HOPS-124 58.29 1.43 41.6 17.8 0.48 2.4E+23 9.9E+06 1.51 / 2.00 438 119.3 80.5
HOPS-182 71.12 1.43 43.7 19.7 0.42 2.1E+23 8.7E+06 1.15 44.8$\pm$4.5 2.25 238 69.7 32.7
HOPS-303 1.49 0.42 20.6 17.3 0.37 1.9E+23 7.8E+06 0.73 / 1.00 207 82.9 53.9
HOPS-361N 478.99 1.70 70.4 23.3 0.36 1.8E+23 7.5E+06 0.92 41.3$\pm$3.2 2.75 202 173.6 45.1
HOPS-361S 478.99 1.10 70.4 23.3 0.23 1.2E+23 4.8E+06 0.88 43.1$\pm$3.8 4.00 139 96.1 40.2
HOPS-384 1477.95 1.98 93.3 33.9 0.28 1.4E+23 5.8E+06 1.02 30.8$\pm$2.2 2.50 32 36.5 67.9
HOPS-403aafootnotemark: 4.14 0.89 22.3 14.8 0.77 3.8E+23 1.6E+07 1.12 / 2.00 479 140.4 42.3
HOPS-404 0.95 0.53 20.4 13.3 0.53 2.6E+23 1.1E+07 0.70 / 1.00 389 59.9 32.0
Complex
HOPS-12W 7.31 0.48 24.8 16.5 0.30 1.5E+23 6.1E+06 0.63 / 2.00 20 20.5 75.3
HOPS-88 15.81 0.70 30.0 18.2 0.34 1.7E+23 7.1E+06 0.76 33.5$\pm$4.8 1.00 71 34.0 53.2
HOPS-373E 5.32 0.53 23.2 15.4 0.43 2.2E+23 9.0E+06 0.68 / 1.50 79 69.8 /
HOPS-373W 5.32 0.64 23.2 15.4 0.52 2.6E+23 1.1E+07 0.71 / 1.50 70 36.0 /
HOPS-398 1.01 0.62 20.4 17.2 0.55 2.8E+23 1.2E+07 0.47 / 1.00 272 76.9 /
HOPS-399 6.34 1.79 24.0 17.2 1.28 6.3E+23 2.6E+07 0.70 46.4$\pm$4.2 1.00 197 66.2 67.6
Unresolved
HOPS-164 0.58 0.24 20.2 13.4 0.19 9.7E+22 4.0E+06 0.77 / 1.00 134 70.6 /
HOPS-224 2.99 0.36 21.5 13.1 0.34 1.7E+23 7.1E+06 0.78 / 1.00 237 107.7 /
HOPS-250 6.79 0.20 24.3 13.5 0.16 7.8E+22 3.3E+06 0.92 / 1.00 134 117.2 /
HOPS-317N 4.76 0.11 22.9 14.2 0.09 4.5E+22 1.9E+06 0.67 / 1.00 52 34.5 /
HOPS-341 2.07 0.24 21.0 14.7 0.23 1.1E+23 4.8E+06 0.63 / 1.00 28 54.5 /
HOPS-402 0.55 0.51 20.2 13.3 0.51 2.5E+23 1.1E+07 0.66 / 1.00 622 156.6 /
HOPS-408 0.52 0.31 20.2 12.2 0.27 1.3E+23 5.6E+06 0.94 / 1.00 125 93.5 /

### v2

# The protostars in Orion: Characterizing the properties of their magnetized envelopes

@@ Abstract
We present a study connecting the physical properties of protostellar envelopes to the morphology of the envelope-scale magnetic field. We used the ALMA polarization observations of 61 young prtostars at 0.87 mm on $\sim 400-3000$ au scales from the B-field Orion Protostellar Survey to infer the envelope-scale magnetic field, and used the dust emission to measure the envelope properties on comparable scales. We find that protostars showing standard-hourglass-field morphology tend to have larger masses and lower velocity dispersions in their envelopes, whereas systems with spiral-field morphologies have higher velocity dispersion. Combining with the disk properties taken from the Orion VLA/ALMA Nascent Disk and Multiplicity survey, we connect envelope properties to fragmentation. Our results show that the fragmentation level is positively correlated with the angle dispersion of the magnetic field, suggesting that the envelope fragmentation tends to be suppressed by the magnetic field. We also find that protostars exhibiting standard hourglass magnetic field structure have a smaller disk and smaller angle dispersion of the magnetic field than other field configurations, specially the rotated hourglass, but also the spiral and others, suggesting a more effective magnetic braking in the standard hourglass morphology of magnetic fields. Nevertheless, significant misalignment between the magnetic field and outflow axes tends to reduce magnetic braking, leading to the formation of larger disks.
@@ 1 Introduction
Fragmentation is believed to be a critical process in the formation of star and stellar systems (e.g., Matsumoto et al., 2003). Initially, turbulence induces chaotic motions and density fluctuations across various scales, compressing the gas in certain regions and leading to localized density enhancements (e.g., Mac Low et al., 2004). Within these dense regions, gravitational forces overcome thermal pressure if the mass exceeds the Jeans mass. As the cloud collapses, the large structures fragment into smaller, dense cores, each of which forms individual stars or multiple star systems (e.g., McKee et al., 2007). This hierarchical fragmentation continues until individual cores reach densities high enough to resist further fragmentation (e.g., Beuther et al., 2015). However, the specific mechanisms by which a cloud core fragments to ultimately form a multiple system or small cluster remain uncertain.
@@
Theoretical and numerical studies (e.g., Maury et al., 2022; Offner et al., 2023) suggest that cores with some angular momentum are prone to fragmentation into multiple objects, indicating the need for additional mechanisms to reduce the fragmentation to the observed levels. Since non-ideal MHD effects do not inhibit protostellar multiplicity (e.g., Wurster et al., 2019), further research points to turbulence and magnetic fields (hereafter, B-fields) as playing a pivotal role in the fragmentation process (e.g., Padoan et al., 2001; Price et al., 2009). For instance, compressive turbulence promotes fragmentation and the formation of structures with high density contrasts, whereas solenoidal turbulence (e.g., Federrath et al., 2010; Girichidis et al., 2011) suppresses fragmentation and maintains a smoother density structure within the cloud. On the other hand, the ratio of rotational to magnetic energy has been found to influence core fragmentation (e.g., Machida et al., 2008). Moreover, some studies suggest that B-field tends to suppress fragmentation by providing additional support against gravitational collapse, and/or through magnetic braking, which carries away the angular momentum from the inner regions of the core (e.g., Commerçon et al., 2011; Guszejnov et al., 2021; Mignon-Risse et al., 2021).
@@
Dust polarization is the most commonly used tracer of B-fields in star-forming regions, as “radiative torques” (e.g., Hoang et al., 2009; Andersson et al., 2015) tend to align spinning and elongated dust grains with their long axes perpendicular to the ambient B-field direction. Over the last decades, dust polarization observations carried out with (sub)millimeter interferometers have increasingly proven effective in mapping B-fields at the scales of cores ($\sim 10^{4}$ au) and envelopes ($\sim 10^{2}$ to $10^{4}$ au) (e.g. Girart et al., 1999, 2013; Cox et al., 2018; Galametz et al., 2018; Hull et al., 2019; Le Gouellec et al., 2020; Cortés et al., 2021). The theory predicts that the B-field regulates the collapse of clouds, whereas ambipolar diffusion allows the formation of “supercritical” dense cores, where the gravitational force dominates over the magnetic support. As cloud cores collapse, B-field lines are dragged inward to be drawn into pinched morphology. This has been observed in various studies (e.g., Girart et al., 2006, 2009; Qiu et al., 2014; Beltrán et al., 2019; Le Gouellec et al., 2019; Hull et al., 2020). In the case of flux freezing, the pinched geometry in the pseudo-disk strengthens the B-field and carries away angular momentum, leading to catastrophic magnetic braking and potentially preventing the disk formation (e.g., Galli et al., 2006). However, disk may form if the magnetic and rotational axes are misaligned (e.g., Joos et al., 2012; Li et al., 2013; Machida et al., 2020). To statistically test these scenarios related to fragmentation and star/disk formation, it is crucial to understand the connection between envelope properties and the resulting fragmentation.
@@
Recently, the first results from the B-fields Orion protostellar survey (BOPS, with a resolution of $\sim 1^{\prime\prime}$, corresponding to $\sim$ 400 au, Huang et al., 2024) presented the largest polarization study of B-fields in protostellar envelopes using dust polarization observations of 61 protostars in Orion. We have successfully mapped the B-fields morphology at scales ranging from several hundred to a few thousand au, and classified the protostars using two criteria: polarization/B-field structure and the alignment between the outflow directions and velocity gradients. On one hand, the morphological categories of the B-field for protostars are classified as: “standard hourglass” (with the main axis coincident with the outflow axis), “rotated hourglass” (similar to the “standard hourglass” but rotated by $\sim 90^{\circ}$, with its axis perpendicular to the outflow), “spiral” configuration, “complex” field morphology, “not enough data” (insufficient polarized data to identify the B-field structure), and “self-scattering” (compact polarized emission in the innermost regions, with the polarization direction perpendicular to the disk major axis). On the other hand, BOPS protostars are classified as “Perp-Type” (velocity gradient is perpendicular to the outflow), “Nonperp-Type” (velocity gradient is not perpendicular to the outflow), and “Unres-Type” (velocity gradient is unresolved). In addition, observations from a major Orion interferometric study, known as the Orion VLA/ALMA Nascent Disk and Multiplicity (VANDAM, Tobin et al., 2020), have identified protostellar disks in the Orion molecular clouds with a resolution of $\sim 0\farcs 1$ (corresponding to $\sim$ 40 au). This study provided data on the radius and mass of these disks. Since all BOPS protostars were originally observed at high resolution in the VANDAM survey, the combination of these datasets enables a comprehensive test of the magnetic braking model and allows for a more effective linkage of magnetic and turbulent properties to the characteristics of protostellar disks and multiplicity.
@@
This paper builds upon the first results of BOPS (Huang et al., 2024) and is organized as follows: Section 2 introduces the observations and data reduction processes. Section 3 presents the methods for the calculation and estimation of physical parameters, the velocity dispersion, the angle dispersion of the B-field, and the fragmentation level. A detailed discussion of the correlation between these parameters is provided in Section 4. Finally, the conclusions are summarized in Section 5.
@@ 2 Observations
The 870 $\mu$m (Band 7) dust polarization observations were taken with ALMA, and the spatial resolution and the Largest Angular Scale (LAS) are $\sim$1″and $\sim$8″, corresponding to a linear resolution of $\sim$400 and $\sim$3200 au at a distance of $\sim$400 pc (see Huang et al. 2024 for further details on the observation setup).
@@ 3 Analysis
The first results of the BOPS have revealed the dust polarized continuum emission for 61 protostars (Huang et al., 2024), and some parameters are listed in Appendix A. In this work, we use the classification criteria based on the B-field/polarization structure because the velocity gradients in a large fraction of the sample are unresolved. Then the 61 protostars are classified into 5 types: “standard hourglass”, “rotated hourglass”, “spiral”, “others” (including “complex” and “not enough data”), and “self-scattering”, as shown in Appendix B. The following sections presents the estimation of physical envelope parameters, velocity dispersion, the angle dispersion of B-field, and fragmentation level for all BOPS protostars. These properties will then be compared to the disk properties to understand which factors could potentially influence the envelope fragmentation and whether the initial misalignment between the B-field and outflows could result in reduced magnetic braking during disk formation.
@@ 3.1 Physical Envelope Properties
where $d$ is the distance, $\nu$ is the frequency corresponding to the observed wavelength, $B_{\nu}(T_{\rm dust})$ is the Planck function at the dust temperature of $T_{\rm dust}$, and $\kappa_{\nu}$ is the dust opacity. We adopt $\kappa_{870\mu{\rm m}}\approx 1.84\ {\rm cm}^{2}\ {\rm g}^{-1}$, which is appropriate for dusty particles with ice mantles at densities of approximately $10^{6}$ cm-3, as reported by Ossenkopf et al. (1994).
@@
Name $L_{\rm bol}$ $S_{\nu}$ $\overline{T}_{\rm env}$ $T_{\rm c}$ $M_{\rm env}$ $N_{\rm gas}$ $n_{\rm gas}$ $\sigma_{\rm nth}$ $\delta\phi$ $N_{F}$ $N_{B}$
@@
HOPS-11 9.00 0.40 26.1 17.7 $\pm$ 0.9 0.26 1.3E+23 5.4E+06 0.67 $\pm$ 0.20 / 1.00 6
HOPS-87 36.49 3.89 37.0 21.2 $\pm$ 1.9 1.46 7.3E+23 3.0E+07 0.43 $\pm$ 0.13 17.6 $\pm$ 1.4 1.00 86
HOPS-359 10.00 0.91 26.8 17.4 $\pm$ 2.0 0.62 3.1E+23 1.3E+07 0.45 $\pm$ 0.11 29.2 $\pm$ 2.3 1.00 80
HOPS-395 0.50 0.58 20.2 14.3 $\pm$ 1.1 0.49 2.4E+23 1.0E+07 0.40 $\pm$ 0.12 15.0 $\pm$ 2.0 2.00 29
HOPS-400 2.94 1.18 21.5 15.8 $\pm$ 0.7 1.01 5.0E+23 2.1E+07 0.45 $\pm$ 0.22 18.2 $\pm$ 1.9 2.00 48
HOPS-407 0.71 0.61 20.2 17.8 $\pm$ 2.0 0.59 2.9E+23 1.2E+07 0.44 $\pm$ 0.12 10.5 $\pm$ 1.0 1.50 53
OMC1N-8-N / 0.63 / 37.5 $\pm$ 10.5 0.31 1.6E+23 6.5E+06 0.43 $\pm$ 0.31 20.6 $\pm$ 2.1 2.00 51
@@
HH270IRS 7.70 0.49 25.1 16.7 $\pm$ 0.1 0.35 1.7E+23 7.2E+06 0.45 $\pm$ 0.15 33.8 $\pm$ 4.1 2.00 35
HOPS-78 8.93 0.81 26.0 22.6 $\pm$ 2.1 0.48 2.4E+23 9.9E+06 0.59 $\pm$ 0.27 40.2 $\pm$ 4.3 2.25 45
HOPS-168 48.07 0.51 39.6 16.5 $\pm$ 1.4 0.17 8.4E+22 3.5E+06 0.64 $\pm$ 0.26 28.4 $\pm$ 3.9 2.25 28
HOPS-169 3.91 0.67 22.2 19.2 $\pm$ 1.4 0.47 2.3E+23 9.7E+06 0.30 $\pm$ 0.18 32.9 $\pm$ 3.9 1.00 37
HOPS-288 135.47 1.60 51.4 13.7 $\pm$ 1.0 0.43 2.1E+23 8.9E+06 1.34 $\pm$ 0.51 24.5 $\pm$ 2.9 3.00 37
HOPS-310 13.83 0.58 29.1 17.0 $\pm$ 1.2 0.33 1.6E+23 6.8E+06 0.75 $\pm$ 0.15 / 1.00 9
HOPS-317S 4.90 2.33 22.9 14.7 $\pm$ 0.9 1.95 9.7E+23 4.0E+07 0.45 $\pm$ 0.16 / 1.25 15
HOPS-370 360.86 1.08 65.6 25.1 $\pm$ 1.9 0.21 1.0E+23 4.3E+06 0.80 $\pm$ 0.32 33.8 $\pm$ 3.1 1.00 62
HOPS-401 0.61 0.37 20.2 14.8 $\pm$ 1.0 0.37 1.9E+23 7.7E+06 0.69 $\pm$ 0.12 / 1.00 13
HOPS-409 8.18 0.38 25.4 30.8 $\pm$ 7.9 0.23 1.2E+23 4.8E+06 0.83 $\pm$ 0.34 17.0 $\pm$ 2.5 1.25 25
OMC1N-4-5-ES / 0.72 / 36.1 $\pm$ 8.9 0.28 1.4E+23 5.8E+06 0.63 $\pm$ 0.21 34.6 $\pm$ 4.1 2.25 66
@@
HOPS-96 6.19 0.44 23.8 25.8 $\pm$ 4.3 0.29 1.4E+23 6.0E+06 0.86 $\pm$ 0.15 / 1.00 11
HOPS-124 58.29 1.43 41.6 20.3 $\pm$ 0.6 0.48 2.4E+23 9.9E+06 1.20 $\pm$ 0.43 / 2.25 8
HOPS-182 71.12 1.43 43.7 16.9 $\pm$ 1.2 0.42 2.1E+23 8.7E+06 1.44 $\pm$ 0.54 46.4 $\pm$ 4.6 3.50 51
HOPS-303 1.49 0.42 20.6 21.4 $\pm$ 1.0 0.37 1.9E+23 7.8E+06 0.56 $\pm$ 0.21 / 1.00 13
HOPS-361N 478.99 1.70 70.4 18.6 $\pm$ 3.1 0.36 1.8E+23 7.5E+06 0.69 $\pm$ 0.33 44.2 $\pm$ 3.4 3.25 86
HOPS-361S 478.99 1.10 70.4 18.8 $\pm$ 3.2 0.23 1.2E+23 4.8E+06 0.66 $\pm$ 0.23 43.1 $\pm$ 3.8 3.50 65
HOPS-384 1477.95 1.98 93.3 51.2 $\pm$ 15.6 0.28 1.4E+23 5.8E+06 0.50 $\pm$ 0.33 30.8 $\pm$ 2.2 2.50 98
HOPS-403 4.14 0.89 22.3 19.0 $\pm$ 3.0 0.77 3.8E+23 1.6E+07 0.80 $\pm$ 0.03 / 2.00 10
HOPS-404 0.95 0.53 20.4 16.0 $\pm$ 0.7 0.53 2.6E+23 1.1E+07 0.37 $\pm$ 0.23 / 1.00 11
Others
HOPS-12W 7.31 0.48 24.8 18.2 $\pm$ 1.4 0.30 1.5E+23 6.1E+06 0.57 $\pm$ 0.18 39.7 $\pm$ 5.0 2.00 23
HOPS-87S / 0.07 / 21.2 $\pm$ 1.8 0.06 2.8E+22 1.2E+06 0.76 $\pm$ 0.18 / 1.00 0
HOPS-88 15.81 0.70 30.0 21.3 $\pm$ 2.1 0.34 1.7E+23 7.1E+06 0.83 $\pm$ 0.25 35.2 $\pm$ 4.3 1.00 35
HOPS-164 0.58 0.24 20.2 15.9 $\pm$ 1.2 0.19 9.7E+22 4.0E+06 0.63 $\pm$ 0.25 / 1.00 7
HOPS-224 2.99 0.36 21.5 13.6 $\pm$ 0.5 0.34 1.7E+23 7.1E+06 0.59 $\pm$ 0.26 / 1.00 6
HOPS-250 6.79 0.20 24.3 14.4 $\pm$ 0.9 0.16 7.8E+22 3.3E+06 0.70 $\pm$ 0.24 / 1.00 5
HOPS-317N 4.76 0.11 22.9 14.7 $\pm$ 0.8 0.09 4.5E+22 1.9E+06 0.37 $\pm$ 0.19 / 1.25 5
HOPS-325 6.20 0.28 23.9 24.6 $\pm$ 5.0 0.22 1.1E+23 4.6E+06 0.44 $\pm$ 0.04 / 1.00 12
HOPS-341 2.07 0.24 21.0 16.0 $\pm$ 0.8 0.23 1.1E+23 4.8E+06 0.19 $\pm$ 0.12 / 1.25 12
HOPS-373E 5.32 0.53 23.2 17.8 $\pm$ 1.5 0.43 2.2E+23 9.0E+06 0.28 $\pm$ 0.18 28.9 $\pm$ 3.6 1.50 33
HOPS-373W 5.32 0.64 23.2 17.8 $\pm$ 1.5 0.52 2.6E+23 1.1E+07 0.50 $\pm$ 0.16 / 1.50 15
HOPS-398 1.01 0.62 20.4 19.8 $\pm$ 2.2 0.55 2.8E+23 1.2E+07 0.50 $\pm$ 0.19 32.0 $\pm$ 5.0 1.00 20
HOPS-399 6.34 1.79 24.0 20.8 $\pm$ 2.3 1.28 6.3E+23 2.6E+07 0.42 $\pm$ 0.15 46.4 $\pm$ 3.2 1.00 103
HOPS-402 0.55 0.51 20.2 14.9 $\pm$ 1.1 0.51 2.5E+23 1.1E+07 0.66 $\pm$ 0.17 / 1.00 10
HOPS-408 0.52 0.31 20.2 13.9 $\pm$ 0.6 0.27 1.3E+23 5.6E+06 / / 1.00 8
OMC1N-4-5-EN / 0.47 / 36.0 $\pm$ 8.7 0.18 9.0E+22 3.8E+06 0.65 $\pm$ 0.26 16.4 $\pm$ 1.4 1.50 66
OMC1N-6-7 / 0.95 / 35.5 $\pm$ 8.8 0.38 1.9E+23 7.8E+06 0.80 $\pm$ 0.58 39.2 $\pm$ 3.8 1.25 53
OMC1N-8-S / 0.16 / 37.0 $\pm$ 10.3 0.06 3.0E+22 1.3E+06 0.48 $\pm$ 0.21 / 1.50 0